How much of an 85% solution of alcohol should be added to 5L of a 45% solution to produce a 60% solution?
Answer by alicealc(293)

let 85% solution of alcohol = x liters

85/100 * x + 45/100 * 5 = 60/100 * (x + 5)
0.85x + 2.25 = 0.6*(x + 5)
0.85x + 2.25 = 0.6x + 3
0.85x - 0.6x = 3 - 2.25
0.25x = 0.75
x = 0.75/0.25 = 75/25 = 3

so we should add 3 liters 85% solution of alcohol